OpenStudy (anonymous):

Use the geometric mean to find the 7th term in a geometric sequence if the 6th term is 75 and the 8th term is 48

OpenStudy (anonymous):

the geoMean of a and b is defined to be: sqrt(ab)

OpenStudy (anonymous):

lets say we have 3 terms of a geometric sequence: n/r, n, nr \[\sqrt{\frac nr*nr}=\sqrt{n^2}=n\]